# AI Chatbots as Professional Service Agents: Developing a Professional Identity

### v1

@@ 5.1 Experimental Settings
Models Approach Flu. Coh. Nat. Emp. Pat-C. Rec. Pre. Llama-2-7B-Chat Zero-Shot 3.90 3.60 2.95 2.10 2.45 0.39 0.07 Zero-Shot + Pragmatic rules 4.00 3.50 3.40 2.70 3.00 0.33 0.10 RAG-based Few-Shot 3.60 2.50 2.65 2.20 2.45 0.38 0.07 Chain-of-Thought 4.00 3.60 2.80 2.20 2.40 0.40 0.07 LAPI 4.00 3.60 3.05 2.90 3.10 0.35 0.07 Llama-2-13B-Chat Zero-Shot 3.75 3.60 2.85 2.25 2.65 0.38 0.08 Zero-Shot + Pragmatic rules 4.00 3.75 3.10 3.05 3.25 0.33 0.12 RAG-Based Few-Shot 3.60 2.65 2.75 2.35 2.25 0.31 0.07 Chain-of-Thought 3.65 3.65 2.95 2.10 2.50 0.39 0.07 LAPI 4.00 3.75 3.15 3.40 3.45 0.36 0.07 Llama-2-70B-Chat Zero-Shot 3.90 3.45 3.15 2.15 2.80 0.35 0.09 Zero-Shot + Pragmatic rules 4.00 3.75 3.40 2.80 3.35 0.35 0.08 RAG-Based Few-Shot 3.80 3.20 2.95 2.35 2.95 0.31 0.11 Chain-of-Thought 3.95 3.45 3.05 2.40 2.70 0.38 0.07 LAPI 4.00 3.75 3.50 3.35 3.35 0.36 0.07 GPT-3.5-Turbo Zero-Shot 4.00 3.60 3.30 2.35 3.15 0.25 0.15 Zero-Shot + Pragmatic rules 4.00 3.60 3.45 3.15 3.55 0.26 0.16 RAG-Based Few-Shot 4.00 3.45 3.25 2.10 3.30 0.26 0.24 Chain-of-Thought 4.00 3.70 3.10 2.30 3.10 0.29 0.16 Semantic Entropy 4.00 3.40 3.10 2.30 3.15 0.21 0.21 LAPI 4.00 3.85 3.45 3.50 3.55 0.30 0.13 GPT-4 Zero-Shot 4.00 3.70 3.05 2.10 2.90 0.35 0.08 Zero-Shot + Pragmatic rules 4.00 3.90 3.50 3.55 3.80 0.35 0.08 RAG-Based Few-Shot 4.00 3.65 3.25 2.30 3.25 0.36 0.11 Chain-of-Thought 4.00 3.70 3.10 2.35 2.85 0.37 0.08 Semantic Entropy 4.00 3.65 3.05 2.40 2.80 0.35 0.07 LAPI 4.00 3.90 3.95 3.85 3.80 0.38 0.07
Table 1: Evaluation results. The table presents the performance of different prompting approaches across various evaluation metrics. Flu., Coh., Nat., Emp., and Pat-C. represent Fluency, Coherence, Naturalness, Empathy, and Patient-Centricity, respectively. Rec. and Pre. refer to ROUGE-L Recall and Precision. Higher scores for the first six metrics (Flu., Coh., Nat., Emp., Pat-C., and Rec.) indicate better performance, while a lower score for Pre. reflects a more desirable outcome. The LAPI approach is often highlighted for its consistency and superior performance across most models and metrics. Bold: Best-performing paradigm per metric.
@@
To comprehensively evaluate the performance of our proposed method, we utilized two categories of metrics: LLM-based and non-LLM-based metrics. For LLM-based evaluation, we employed G-EVAL Liu et al. (2023), a state-of-the-art metric designed to assess various qualitative aspects of large language model outputs, including Fluency, Coherence, Naturalness, Empathy, and Patient-Centricity, where higher scores indicate better performance. Detailed descriptions of these five evaluation metrics can be found in Appendix B. G-EVAL has been shown to exhibit strong consistency with human evaluations Zhu et al. (2025); Gao et al. (2024), making it a reliable tool for qualitative assessment. For non-LLM-based evaluation, we adopted ROUGE-L recall and precision Lin (2004) as objective n-gram matching-based metrics. ROUGE-L Recall measures the extent to which the generated responses cover the content of the doctor-provided answers, with higher values indicating better coverage, while ROUGE-L Precision evaluates the informativeness of the generated responses, with higher values indicating that the generated responses contain more information.
Figure 4: Overview of GPT-3.5-Turbo’s responses using different methods to answer the patient query: ”After undergoing laser surgery for high myopia, will the degree of myopia increase again?” The text in green provides the direct answer to the question, the text in yellow encourages regular contact with the patient’s doctor and the text in blue offers actionable advice for the patient.
@@
We categorize the baseline methods into two groups: prompt design-based methods and quality inspection-based methods. For the prompt design-based methods, we explored strategies to guide the model’s reasoning and response generation. Specifically, we include the following: Zero-Shot, where the model generates responses directly without additional guidance; Zero-Shot + Pragmatic Rules, which enhances the zero-shot approach by incorporating explicit pragmatic rules tailored to professional contexts; RAG-based Few-Shot Ori et al. (2023); Guoxin et al. (2023), a retrieval-augmented generation strategy that retrieves relevant examples from a knowledge base to assist response generation; and Chain-of-Thought (CoT) Kojima et al. (2023), which encourages the model to generate step-by-step reasoning paths to improve response accuracy and relevance. For the quality inspection-based methods, we employed Semantic Entropy Farquhar et al. (2024), a technique designed to evaluate the confidence of the model regarding specific knowledge points in its generated responses. The confidence scores are then fed back into the model, prompting it to revise and improve its outputs.
@@
We evaluate our proposed method on five large language models (LLMs): Llama-2-7B-Chat, Llama-2-13B-Chat, Llama-2-70B-Chat Touvron et al. (2023b), GPT-3.5-Turbo (gpt-3.5-turbo-0125), and GPT-4 (gpt-4-0125-preview), selected for their diverse scales and capabilities, representing both open-source and closed-source models. To enhance the alignment of responses with professional and human-like communication, we introduced pragmatic rules used in both Zero-Shot+Pragmatic Rules and LAPI approaches, as detailed in Appendix C.
For the RAG-based Few-Shot method, we employed a retrieval-augmented generation strategy using a BERT encoder to retrieve the three most relevant question-answer pairs based on cosine similarity, which were then used as prompts to guide response generation. In the HBM + Pragmatic Entropy method, we set $n$ to 5, $\tau$ to 0.5, and $\epsilon$ to 0.8. For the Semantic Entropy method, due to the high text comprehension capability required for this method, Semantic Entropy was only applied to GPT-3.5-Turbo and GPT-4.
@@ 5.4 Case Study
To further illustrate the capabilities of our proposed method, we present a case study using GPT-3.5-Turbo to answer the patient query: ”After undergoing laser surgery for high myopia, will the degree of myopia increase again?”. Figure 4 provides a comparison of the responses generated by four different prompting strategies: LAPI, Chain-of-Thought, Zero-Shot, and RAG-Based Few-Shot. In the responses, the text in green represents the direct answer to the question, the text in yellow encourages regular contact with the patient’s doctor, and the text in blue offers actionable advice for the patient.
As observed in the case study, LAPI (left) outperforms methods like Chain-of-Thought, Zero-Shot, and RAG-Based Few-Shot. LAPI gives a more comprehensive and detailed response, directly answering the question, stressing the importance of regular ophthalmologist follow-ups, and offering advice on healthy eye habits. This showcases the benefit of integrating the HBM framework and Pragmatic Entropy. In contrast, other methods provide direct answers and some patient advice but lack LAPI’s level of detail and actionable guidance. For instance, Chain-of-Thought misses lifestyle advice, and RAG-Based Few-Shot doesn’t emphasize continuous patient-doctor communication as strongly as LAPI.
@@ 6 Conclusion
In this work, we introduce a novel approach for designing LLM-based professional service agents tailored for medical Q&A services. Our method combines theory-guided task planning and pragmatic entropy to ensure professional-aligned communication for ethical, patient-centric responses. Experiments show that our method outperforms baseline techniques in fluency, empathy, patient-centricity, and ROUGE-L scores. This research highlights the potential of enhancing AI chatbot communication, particularly in healthcare, and encourages further development of more effective methods in future work.

### v2

@@ 5.1 Experimental Settings
Models Approach Flu. Coh. Nat. Emp. Pat-C. Rec. Pre. Llama-2-7B-Chat Zero-Shot 3.90 3.60 2.95 2.10 2.45 0.39 0.07 Zero-Shot + Pragmatic rules 4.00 3.50 3.40 2.70 3.00 0.33 0.10 RAG-based Few-Shot 3.60 2.50 2.65 2.20 2.45 0.38 0.07 Chain-of-Thought 4.00 3.60 2.80 2.20 2.40 0.40 0.07 LAPI 4.00 3.60 3.05 2.90 3.10 0.35 0.07 Llama-2-13B-Chat Zero-Shot 3.75 3.60 2.85 2.25 2.65 0.38 0.08 Zero-Shot + Pragmatic rules 4.00 3.75 3.10 3.05 3.25 0.33 0.12 RAG-Based Few-Shot 3.60 2.65 2.75 2.35 2.25 0.31 0.07 Chain-of-Thought 3.65 3.65 2.95 2.10 2.50 0.39 0.07 LAPI 4.00 3.75 3.15 3.40 3.45 0.36 0.07 Llama-2-70B-Chat Zero-Shot 3.90 3.45 3.15 2.15 2.80 0.35 0.09 Zero-Shot + Pragmatic rules 4.00 3.75 3.40 2.80 3.35 0.35 0.08 RAG-Based Few-Shot 3.80 3.20 2.95 2.35 2.95 0.31 0.11 Chain-of-Thought 3.95 3.45 3.05 2.40 2.70 0.38 0.07 LAPI 4.00 3.75 3.50 3.35 3.35 0.36 0.07 GPT-3.5-Turbo Zero-Shot 4.00 3.60 3.30 2.35 3.15 0.25 0.15 Zero-Shot + Pragmatic rules 4.00 3.60 3.45 3.15 3.55 0.26 0.16 RAG-Based Few-Shot 4.00 3.45 3.25 2.10 3.30 0.26 0.24 Chain-of-Thought 4.00 3.70 3.10 2.30 3.10 0.29 0.16 Semantic Entropy 4.00 3.40 3.10 2.30 3.15 0.21 0.21 LAPI 4.00 3.85 3.60 3.50 3.70 0.29 0.13 GPT-4 Zero-Shot 4.00 3.70 3.05 2.10 2.90 0.35 0.08 Zero-Shot + Pragmatic rules 4.00 3.90 3.50 3.55 3.80 0.35 0.08 RAG-Based Few-Shot 4.00 3.65 3.25 2.30 3.25 0.36 0.11 Chain-of-Thought 4.00 3.70 3.10 2.35 2.85 0.37 0.08 Semantic Entropy 4.00 3.65 3.05 2.40 2.80 0.35 0.07 LAPI 4.00 3.90 3.95 3.85 3.80 0.38 0.07
Table 1: Evaluation results. The table presents the performance of different prompting approaches across various evaluation metrics. Flu., Coh., Nat., Emp., and Pat-C. represent Fluency, Coherence, Naturalness, Empathy, and Patient-Centricity, respectively. Rec. and Pre. refer to ROUGE-L Recall and Precision. Higher scores for the first six metrics indicate better performance, while a lower score for Pre. reflects a more desirable outcome. Bold: Best-performing paradigm per metric.
Figure 3: Overview of GPT-3.5-Turbo’s responses using different methods to answer the patient query: "After undergoing laser surgery for high myopia, will the degree of myopia increase again?"
@@
To comprehensively evaluate the performance of our proposed method, we utilized two categories of metrics: LLM-based and non-LLM-based metrics. For LLM-based evaluation, we employed G-EVAL[liu-etal-2023-g], a state-of-the-art metric designed to assess various qualitative aspects of large language model outputs, including Fluency, Coherence, Naturalness, Empathy, and Patient-Centricity, where higher scores indicate better performance. Detailed descriptions of these five evaluation met- rics can be found in Appendix B. G-EVAL has been shown to exhibit strong consistency with human evaluations[zhu-etal-2025-factual, gao2024largelanguagemodelswikipediastyle], making it a reliable tool for qualitative assessment. For non-LLM-based evaluation, we adopted ROUGE-L recall and precision[lin-2004-rouge] as objective n-gram matching-based metrics. ROUGE-L Recall measures the extent to which the generated responses cover the content of the doctor-provided answers, with higher values indicating better coverage, while ROUGE-L Precision evaluates the informativeness of the generated responses, with lower values indicating that the generated responses contain more information.
@@
We categorize the baseline methods into two groups: prompt design-based methods and quality inspection-based methods. For the prompt design-based methods, we include the following: Zero-Shot, where the model generates responses directly without additional guidance; Zero-Shot + Pragmatic Rules, which enhances the zero-shot approach by incorporating explicit pragmatic rules tailored to professional contexts; RAG-based Few-Shot[ram-etal-2023-context, yu-etal-2023-retrieval], a retrieval-augmented generation strategy that retrieves relevant examples from a knowledge base to assist response generation; and Chain-of-Thought (CoT)[kojima2023largelanguagemodelszeroshot], which encourages the model to generate step-by-step reasoning paths to improve response accuracy and relevance. For the quality inspection-based methods, we employed Semantic Entropy[farquhar-etal:detecting-hallucinations], a technique designed to evaluate the confidence of the model regarding specific knowledge points in its generated responses. The confidence scores are then fed back into the model, prompting it to revise and improve its outputs.
@@
We evaluate our proposed method on five LLMs: Llama-2-7B-Chat, Llama-2-13B-Chat, Llama-2-70B-Chat [touvron2023llama2openfoundation], GPT-3.5-Turbo (gpt-3.5-turbo-0125), and GPT-4 (gpt-4-0125-preview), selected for their diverse scales and capabilities, representing both open-source and closed-source models. To enhance the alignment of responses with professional and human-like communication, we incorporated pragmatic rules developed through a streamlined process: we reviewed medical literature, identified key communication principles (focusing on reply format, information content, and user emotional impact), and distilled these into actionable rules, which are detailed in Appendix C.
For the RAG-based Few-Shot method, we employed a retrieval-augmented generation strategy using a BERT encoder to retrieve the three most relevant question-answer pairs based on cosine similarity, which were then used as prompts to guide response generation. In the LAPI method, we set $n$ to 5, $\tau$ to 0.5, and $\epsilon$ to 0.8. For the Semantic Entropy method, due to the high text comprehension capability required for this method, which was only applied to GPT models.
@@ 5.4 Case Study
To further illustrate the capabilities of our proposed method, we present a case study using GPT-3.5-Turbo to answer the patient query. Figure 3 provides a comparison of the responses generated by four different prompting strategies: LAPI, CoT, Zero-Shot, and RAG-Based Few-Shot. In the responses, green text directly answers the question, yellow text promotes regular doctor contact, and blue text provides actionable patient advice. As observed in the case study, LAPI (left) outperforms others, providing a more comprehensive and detailed response that directly answers the question, stresses the importance of regular ophthalmologist follow-ups, and offers advice on healthy eye habits. This showcases the benefit of integrating the HBM framework and Pragmatic Entropy. In contrast, other methods provide direct answers and some patient advice but lack LAPI’s level of detail and actionable guidance. For instance, CoT misses lifestyle advice, and RAG-Based Few-Shot doesn’t emphasize continuous patient-doctor communication as strongly as LAPI.
The LAPI framework also demonstrates superior performance compared to various CoT variants, with detailed results presented in Appendix D.
@@ 6 Conclusion
In this work, we propose a novel approach for designing LLM-based professional service agents specifically tailored for medical Q&A services. Our method promotes communication aligned with professional identities by integrating theory-guided task planning and a pragmatic entropy method to generate ethical, coherent, and patient-centric responses. Experimental results demonstrate that our approach outperforms baseline methods in fluency, empathy, patient-centricity, and ROUGE-L scores. This study underscores the potential of enhancing AI chatbot communication in healthcare and encourages the continued development of more effective methods in future research.
7 Limitations
While our LAPI framework shows significant promise for enhancing professional communication in medical Q&A services, several aspects require further refinement. The current evaluation primarily relies on large language models to assess output quality, which aligns closely with human judgment. However, incorporating real-world testing, such as deploying the model in clinical settings, would further validate its practical effectiveness and applicability. While this didn’t affect the validity of the proposed method—since results across different LLM sizes showed notable improvements over baseline methods in attributes like Empathy—the smaller sample size might limit how broadly the findings can be applied. Furthermore, the experiments focused solely on medical Q&A scenarios. While the framework is designed to apply to other domains like law, education, or finance, its effectiveness in these areas requires further validation.
8 Ethical statements
The dataset was provided by medical professionals from a reputable ophthalmology hospital and consists solely of anonymized question-answer pairs without any identifiable patient information. During the data collection process, patients provided verbal consent for their questions to be used for academic research purposes and were advised to avoid including personal or identifiable information in their queries. This process ensured that no personal health information was collected or used. Furthermore, the dataset was reviewed and approved by the Institutional Review Board (IRB) of the collaborating ophthalmology hospital prior to its use in our study, ensuring compliance with ethical standards.
9 Acknowledgements
We would like to thank all the reviewers for their helpful feedback, and EMNLP 2025 and ACL Rolling Review organizers for their efforts. This work was supported by the grants from National Natural Science Foundation of China (No.72201068, 72131004, 72342011).
Appendix A Prompts Used in Theory-guided Task Planning
Figure 5: Prompts used to generate an initial prompt that accurately reflects the correct placement of the question within the HBM theory framework.
Appendix B Evaluation Metrics
In the following content, we provide a detailed description of the five key evaluation metrics used to assess the performance of the proposed method, all of which are evaluated using the G-EVAL[liu-etal-2023-g] framework with GPT-4. This approach ensures a comprehensive and consistent evaluation of the generated responses.
  • Fluency (1–4 points): Quality of the response in terms of grammar, spelling, punctuation, word choice, and sentence structure.
    • 4 points: No noticeable issues with grammar, spelling, punctuation, word choice, and sentence structure; the response is concise and clear.
    • 3 points: The response has few or no grammatical or spelling errors; some word choices or structures may seem slightly unnatural but do not impede understanding.
    • 2 points: There are noticeable grammatical, spelling, or structural issues that affect fluency, but the overall message is understandable.
    • 1 point: Many errors in grammar, spelling, punctuation, or word choice, affecting overall understanding and the fluency of the response.
  • Coherence (1–4 points): This dimension aligns with structure and coherence in the DUC quality issues, where the answer should be well-structured and organized. The answer should not merely be a compilation of related information but should develop step-by-step into a coherent body of information on a specific topic.
    • 4 points: Clear structure and logical organization, with natural and smooth transitions between sentences. Not only is the information relevant, but it also demonstrates overall coherence of the response, avoiding harsh transitions or topic interruptions.
    • 3 points: Generally clear structure and organization, though occasional logical jumps or transitions may not be natural. The thematic development is mostly complete, but the connection of some content is slightly mechanical or abrupt, slightly diminishing the overall coherence.
    • 2 points: Clear breaks or jumps between sentences, with rather loose logical organization. The thematic development of the information is unclear, with some content appearing irrelevant or repetitive, making the overall presentation disjointed.
    • 1 point: Sentences lack a clear logical relationship, with chaotic content organization, making it difficult to form a coherent theme development. There is a significant feeling of information piling up, lacking structural organization, making it difficult for the reader to understand the core idea.
  • Naturalness (1–4 points): Assesses whether the answer conforms to the habits of natural language expression, resembling the tone and logic of human everyday communication. Focus on word choice, tone, sentence structure, and the overall reading experience.
    • 4 points: The expression is smooth, the tone is natural and close to everyday conversation, sentence structures are flexible and varied, with no stiffness. Word choice is precise and engaging, creating a sense of realism in the dialogue.
    • 3 points: Overall, the expression is natural, but some parts may appear formulaic or mechanized. The tone is slightly flat but does not hinder understanding. Word choice is appropriate but may lack variety.
    • 2 points: Language expression is partly stiff or awkward, the tone is unnatural, appearing mechanical or overly formal. There may be sentence structures or logical arrangements that do not conform to everyday language habits.
    • 1 point: The mode of expression is clearly unnatural, the tone is stiff or stilted, completely detached from everyday language habits, difficult for people to understand or accept.
  • Empathy (1–4 points): Comprehensive assessment of an individual’s ability to understand, empathize with, and respond to others’ emotions.
    • 4 points: Able to deeply understand and empathize with others’ emotions, responding appropriately to others’ feelings, demonstrating high emotional sensitivity and care.
    • 3 points: Can understand others’ emotions and appropriately respond to most situations, but may occasionally fall short in complex or subtle emotional expressions.
    • 2 points: Sometimes understands others’ emotions, but responses may lack depth or contain misunderstandings, showing some emotional detachment.
    • 1 point: Struggles to understand others’ emotions or emotional reactions, lacks empathy, responses often seem inappropriate or irrelevant.
  • Patient-Centricity (1–4 points): Comprehensive assessment of an individual’s ability to center communication on the patient’s needs, providing clear, practical, and approachable information in a conversational manner.
    • 4 points: Communication is highly patient-centered, using concise and approachable language, delivering practical advice tailored to the patient’s situation, and maintaining a logical and conversational tone that fosters trust and understanding. Without listing ideas.
    • 3 points: Generally patient-centered, but may lack full clarity, practical detail, or conversational ease in some parts while still being largely useful and understandable.
    • 2 points: Moderately patient-centered but may include excessive technical details, lack practical guidance, or have a less approachable tone, making it harder for patients to follow or apply.
    • 1 point: Communication fails to center on the patient, being overly formal, academic, or irrelevant, and lacking clarity, practicality, or approachability.
Appendix C Pragmatic Rules
Our pragmatic rule acquisition followed a rigorous and multi-phase process. First, we conducted a systematic review of medical literature. Second, we summarized the principles that doctors should follow in communication with patients into three aspects: what format to use for replies, what information to convey, and how the user will feel. Third, we distilled actionable rules from these principles. The pragmatic rules used in this paper are listed below.
  • Without listing ideas, reply like a human being[longwen:how].
  • The reply is easy to understand, and the technical terms have relevant explanations[Katherine:jargon].
  • Showing empathy: a comprehensive assessment of an individual’s ability to understand, empathize with, and respond to others’ emotions[xiaoyu:physicianempathy].
  • The reply has authenticity and feasibility[CHEN2020102253].
  • Clear structure and logical organization, with natural and smooth transitions between sentences[JIANG2024108161].
Appendix D LAPI vs. CoT Variants Performance with GPT-4
To demonstrate the distinct advantages of the LAPI framework over Chain-of-Thought (CoT) prompting, particularly in providing actionable advice, we conducted a comparative analysis between LAPI and several enhanced CoT variants. This evaluation aims to verify whether LAPI’s superior performance persists even when CoT prompts are explicitly designed to align more closely with LAPI’s approach, thus clarifying LAPI’s technical contribution. Specifically, we compared LAPI against four CoT variants using GPT-4: (1) Baseline CoT with ”Let’s think step by step” prompting; (2) CoT-(P.R.), which incorporates pragmatic rules; (3) CoT-(HBM), which integrates HBM theory; and (4) CoT-(HBM+P.R.), which adds pragmatic rules to CoT-(HBM). As shown in Table 3, LAPI consistently outperforms all CoT variants across multiple metrics, reinforcing its distinct technical advantages.
Approach Flu. Coh. Nat. Emp. Pat-C. Rec. Pre.
CoT 4.00 3.70 3.10 2.35 2.85 0.37 0.08
CoT-(P.R.) 4.00 3.90 3.75 3.65 3.55 0.36 0.07
CoT-(HBM) 4.00 3.85 3.15 2.80 3.05 0.36 0.08
CoT-(HBM+P.R.) 4.00 3.80 3.90 3.60 3.75 0.35 0.07
LAPI 4.00 3.90 3.95 3.85 3.80 0.38 0.07
Table 3: Performance comparison between LAPI and improved CoT prompts using GPT-4.